# HallE-Switch: Rethinking and Controlling Object Existence Hallucinations in Large Vision Language Models for Detailed Caption

### v1

@@ Abstract
Current large vision-language models (LVLMs) achieve remarkable progress, yet there remains significant uncertainty regarding their ability to accurately apprehend visual details, that is, in performing detailed captioning. To address this, we introduce CCEval, a GPT-4 assisted evaluation method tailored for detailed captioning. Interestingly, while LVLMs demonstrate minimal object existence hallucination in existing VQA benchmarks, our proposed evaluation reveals continued susceptibility to such hallucinations. In this paper, we make the first attempt to investigate and attribute such hallucinations, including image resolution, the language decoder size, and instruction data amount, quality, granularity. Our findings underscore the unwarranted inference when the language description includes details at a finer object granularity than what the vision module can ground or verify, thus inducing hallucination. To control such hallucinations, we further attribute the reliability of captioning to contextual knowledge (involving only contextually grounded objects) and parametric knowledge (containing inferred objects by the model). Thus, we introduce HallE-Switch, a controllable LVLM in terms of Hallucination in object Existence. HallE-Switch can condition the captioning to shift between (i) exclusively depicting contextual knowledge for grounded objects and (ii) blending it with parametric knowledge to imagine inferred objects. Our method reduces hallucination by 44% compared to LLaVA and maintains the same object coverage.
@@ 1 Introduction
In recent years, Large Vision-Language Models (LVLMs) (Liu et al., 2023a; Dai et al., 2023; Li et al., 2023b; Zhu et al., 2023) have achieved significant progress, advancing tasks such as detailed captioning, visual conversations, and vision question-answering (VQA) (Goyal et al., 2017; Liu et al., 2023b; Hudson & Manning, 2019; Fu et al., 2023). However, similar to Large Language Models (LLMs) (Touvron et al., 2023a; Team, 2023; OpenAI, 2022) in the NLP domain, LVLMs confront the issue of hallucination. This challenge is especially severe in detailed image captioning, crucial for accurately describing visual details and supporting applications in robotics (Huang et al., 2023), visual search (Hu et al., 2023), etc. To better understand and address this challenge, we first outline the various types of hallucinations frequently observed in the detailed captions:
Object Existence Hallucination: The detailed image description references objects that are not present.
Object Attribute Hallucination: The detailed image description inaccurately characterizes objects, misrepresenting attributes such as color, shape, and size.
Object Relationship Hallucination: The detailed image description inaccurately depicts the relationships or interactions among objects, including erroneous relative positions, interaction states, and actions involving two or more objects.
@@
Evaluating detailed captions is inherently complex. Some of the efforts, including benchmarks like POPE (Li et al., 2023d), evaluate object hallucination using VQA. Such a bias towards VQA-based evaluations might result in an incomplete assessment of detailed captions, particularly in obtaining a comprehensive view of visual details. To bridge this gap, we introduce CCEval, designed specifically for object existence hallucination in detailed captions. To avoid the model gaining an unfair advantage by favoring shorter descriptions, CCEval maintains consistency in metrics such as average sentence length and the number of objects. Notably, even models that well-performed on VQA-based object hallucination benchmarks showed substantial hallucinations when evaluated with CCEval. Given the finding, we systematically tackle hallucinations, with an initial focus on object existence hallucination.
@@
In our exploration to uncover the underlying causes of object existence hallucination, we find various factors including the size of the language decoder, the quantity, quality, and granularity of instruction data, and the input resolution to the vision encoder. We observe that increasing the language decoder size can reduce hallucination, but more crucially, vision encoder needs to ground the object accurately. As illustrated in Figure 1(a), during training of LVLMs, the goal is to establish a one-to-one correspondence between objects mentioned in the caption and those present in the image. Objects successfully grounded by the vision encoder form accurate associations, internalizing them as contextual knowledge. Conversely, objects in the language that the vision encoder fails to ground create word-word semantic associations, which can be attributed to the generalization from the parametric knowledge within the model’s parameters. During inference, when the model draws from such parametric knowledge, any misalignment can manifest as hallucination, as the model attempts to “guess” details not grounded by the vision module.
@@
Recognizing the significance of both contextual and parametric knowledge in ensuring generation reliability, we present HallE-Switch, a novel approach to control the extent of expressed hallucination or parametric knowledge. The motivation is that not all hallucination is bad, and it is more desirable to control the generalization rather than outright removal of all imagined objects. We curate a 33k dataset similar to LLaVA (Liu et al., 2023a), incorporating both pure contextual knowledge and a blend of contextual knowledge with marked parametric knowledge. Leveraging this dataset, we train a lightweighted single linear layer to control over the frozen LVLM. As demonstrated in Figure 1(b), a singular continuous parameter adjustment (e.g. ) during inference enables the model to produce detailed captions with only contextual knowledge (e.g., ) or blend with parametric knowledge (e.g., ). Furthermore, the inferred objects from parametric knowledge are automatically highlighted with distinct tokens (e.g., [object]) for human reference. This method offers the advantage of preserving object count and coverage as well as sentence length, while effectively control object existence hallucination.
@@
A novel evaluation method for detailed caption object existence hallucination, with metrics such as object count, coverage, and average sentence length, alongside hallucination assessment;
A comprehensive analysis on LVLM components that influence hallucination, with a specific focus on alignment issues in the vision encoder and instruction data;
An approach to control object existence hallucination within detailed captions.
@@ 2 Hallucination Analysis
Object existence hallucination can be influenced by several factors, including the language decoder, instruction data, and vision encoder. In our analysis, we address each factor individually. For a diverse methodological exploration, we select LLaVA, InstructBLIP (Dai et al., 2023), and Shikra (Chen et al., 2023): LLaVA and Shikra share the same model structure; Shikra and InstructBLIP use mixed-dataset and multi-task instruction data; InstructBLIP finetunes only Q-former, while the other finetune projector and LLM.
@@ 2.1 Benchmarks
There are two primary approach, VQA-based and caption-based benchmarks, for evaluating object existence hallucination in LVLMs.
@@
VQA-based benchmarks pose questions about details of objects within images. For a model to be considered hallucination-free, it should address these visual questions accurately. Notably, a large proportion of these questions are simply binary, typically asking about the presence or attributes of objects.
The POPE benchmark evaluates object existence hallucination by a polling-based query method, consisting of a series of yes/no questions on sampled objects from visual instructions. POPE contains three sets: random, popular, and adversarial. These subsets respectively focus on randomly selected objects, frequently occurring objects, and those objects that co-occur in training sets. We choose POPE evaluation on MSCOCO (Lin et al., 2014) dataset. MME coarse-grained recognition construct yes/no questions similarly but selects objects at random. This benchmark has 30 images, with each image paired with two questions: one positive and one negative.
In Table 1, LLaVA exhibits the greatest degree of hallucination, whereas Shikra outperforms other models in both POPE and MME. Specifically, Shikra shows a significantly higher F1 score in both POPE-popular and POPE-adversarial categories, while LLaVA displays the lowest. Additionally, Shikra’s ”Yes” ratio is closer to a balanced 50% compared to other models. However, in subsequent sections, we demonstrate that these observations from VQA-based benchmarks are not consistent with those from caption-based benchmarks.
@@
| CHAIR | CHAIR | Avg. Length | Avg. Object | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object | |
| LLaVA | 24.1 | 9.1 | 42.5 | 3.7 | 72.00 | 19.7 | 32.74 | 92.27 | 9.19 |
| LLaVA | 60.6 | 18.4 | 90.2 | 7.6 | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 |
| Shikra | 59.1 | 16.6 | 91.2 | 7.5 | 83.00 | 24.40 | 33.29 | 109.37 | 9.10 |
| InstructBLIP | 1.4 | 1.7 | 2.3 | 0.8 | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 |
@@
Table 2(left) reveals that while InstructBLIP exhibits minimal object existence hallucination, it averages a mere 0.8 objects per sentence. In contrast, LLaVA and Shikra manifest a higher degree of hallucination, but they also generate more detailed captions, outputting as many as 7.6 and 7.5 objects per sentence, respectively. We find infeasible to comparing object hallucinations is impractical when there is a significant disparity in average sentence length and the number of objects.
Apart from the disparities in the average number of objects, the use of a hard-coded ground truth set is another challenge. To counter these challenges, we introduce CCEval, a GPT-4 assisted evaluation for detailed captions. We first prompt LVLMs to generate detailed captions. Subsequently, utilizing GPT-4’s in-context learning capabilities, we extract individual objects from these captions and identify hallucinated ones by referencing the provided ground truth objects. On top of CHAIR, we introduce ”coverage” metric to ensure that the captions are detailed enough. This metric computes the ratio of objects in the caption that match the ground truth to the total number of ground truth objects from Visual Genome (Krishna et al., 2017). We additionally record and and balance the average number of objects and the average length of captions across all cases.
As reflected in Table 2(right), when subjected to consistent constraints—average sentence length approximately 100 words and around 9 objects per sentence—all models exhibit comparable results. Interestingly, while Shikra surpass other models in VQA-based benchmarks, especially in the POPE , it under-performed in CCEval. This suggests that object existence hallucination in detailed captions is consistently captured by VQA-based evaluations.
@@
| POPE - Popular | LLaVA | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 |
@@
| POPE - Adversarial | LLaVA | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
@@
| Benchmark | Model | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | LLaVA | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
@@ 2.2 Language Decoder
We investigate if expanding the size of the language backbone can mitigate object existence hallucination. As detailed in Table 3, the language decoder of LLaVA is increased from 7B to 33B, and for InstructBLIP, it is increased from 7B to 13B. The result shows that hallucination for LLaVA reduced for POPE but not for CCEval. For InstructBLIP, and is reduced by 8 and 5.6 on CCEval, respectively. More importantly, although there is a gain for scaling up language backbone, the performance does not scale linearly, as the accuracy increased 5.73 from 7B to 13B and only 2.37 from 13B to 33B on POPE-adversarial.
@@ 2.3 Data
Similar to our approach with the language decoder, we begin by scaling up the volume of instruction finetuning data, ranging from 80K to 2.4M. As illustrated in Table 4, the LLaVA model, finetuned on 80K instruction data, exhibits fewer object existence hallucinations compared to the models finetuned on 150K and SVIT (Zhao et al., 2023). The result suggests extra data may increase hallucination for both VQA-based and caption-based evaluations. Given that all three datasets are generated by GPT-4, we question the quality of the data. LRV also raises this concern, suggesting that the training data itself might contain hallucinations. Some examples of training data are presented in the Appendix. Interestingly, our examination shows no object existence hallucination: the objects in the captions are contained the ground truth objects from MSCOCO. However, we identify certain ground truth objects are challenging for human observers to ground, due to factors like size, resolution, and occlusion. This led us to hypothesize that the vision encoder might also struggle to ground these objects effectively.
@@
| Benchmark | Model | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
@@
| Benchmark | Vision Encoder | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
@@
| Benchmark | Vision Encoder | CHAIR | CHAIR | Coverage | Avg. Length | Avg. Object |
@@ 2.4 Vision Encoder
Intuitively, increasing image resolution enhances model’s perception of finer details, thus making the grounding of objects mentioned in the caption easier. To verify our hypothesis from the previous section, we increase the input image resolution for the vision encoder. Specifically, for our evaluation, the resolution for LLaVA was incremented from 224x to full resolution using a sliding window approach for efficiency, as detailed in the Appendix. Table 6 shows a constant decrease in hallucination and increase in object coverage. Additionally, we assess LLaVA with Llama 2, varying the resolution from 112x to 336x. For the 112x112 resolution, the original image was downscaled to 112x and subsequently upscaled to 224x before being input to CLIP-Large-224x. Table 5 gives a consistent observation that larger input resolution can reduce hallucination.
Conclusion. Through our systematic analysis of object existence hallucination, we summarize several insights: (1) Enlarging the language decoder does mitigate hallucination, but gains diminish while scaling up, indicating it is not the primary factor. (2) Expanding the volume of instruction data actually increases hallucination. Upon inspection of training data, we find certain objects described in the captions might not be grounded by the vision encoder. (3) Improving input image resolution significantly reduces hallucination by enhancing grounding, validating our hypothesis.
Reflecting on these findings, we attempt to provide an explanation for the reason of object existence hallucination in detailed captions. The process of image captioning in LVLMs can be perceived as a form of information mapping or translation. Ideally, the goal is to have a direct one-to-one correspondence between objects identified in the image and those mentioned in the captions. Objects successfully grounded by the vision encoder form accurate associations. We define this kind of knowledge as contextual knowledge in the model, following (Neeman et al., 2023). We define parametric knowledge, the knowledge is encoded in the model’s parameter, to be the association of objects in the language with other words instead of with corresponding image object feature, formed when they fails to ground by the vision encoder. During inference, when the model draws from this parametric knowledge, it attempts to “guess” details not grounded by the vision module and is perceived as object existence hallucination. However, such guessing is necessary depending the details required by the downstream task. Therefore, our approach work towards controlling and balancing parametric knowledge rather than eliminating.

### v3

# HallE-Control: Controlling Object Hallucination in Large Multimodal Models

@@ Abstract
Current Large Multimodal Models (LMMs) achieve remarkable progress, yet there remains significant uncertainty regarding their ability to accurately apprehend visual details, that is, in performing detailed captioning. To address this, we introduce CCEval, a GPT-4 assisted evaluation method for detailed captioning. Interestingly, while LMMs demonstrate minimal object existence hallucination in existing VQA benchmarks, our proposed evaluation reveals continued susceptibility to such hallucinations. In this paper, we make the first attempt to investigate such hallucination from different aspects, including image resolution, the language decoder size, and instruction data amount, quality, granularity. Our findings underscore the unwarranted inference when the language description includes details at a finer object granularity than what the vision module can ground or verify, thus inducing hallucination. To control such hallucinations, we further attribute the reliability of captioning to contextual knowledge (involving only contextually grounded objects) and parametric knowledge (containing inferred objects by the model). Thus, we introduce HallE-Control, a controllable LMM in terms of Hallucination in object Existence. HallE-Control can condition the captioning to shift between (i) exclusively depicting contextual knowledge for grounded objects and (ii) blending it with parametric knowledge to imagine inferred objects. Our method reduces hallucination by 44% compared to LLaVA7B and maintains the object coverage. Our code is publicly available at https://github.com/bronyayang/HallE_Control
@@ 1 Introduction
In recent years, Large Multimodal Models (LMMs) (Liu et al., 2023d; Dai et al., 2023; Li et al., 2023b; Zhu et al., 2023) have achieved significant progress, advancing tasks such as detailed captioning, visual conversations, and vision question-answering (VQA) (Goyal et al., 2017; Liu et al., 2023e; Hudson & Manning, 2019; Fu et al., 2023). However, similar to Large Language Models (LLMs) (Touvron et al., 2023a; Team, 2023; OpenAI, 2022) in the NLP domain, LMMs confront the issue of hallucination. This is particularly severe in detailed image captioning, which hinders the performance of downstream applications in robotics (Huang et al., 2023), visual search (Hu et al., 2023), etc. To better understand and address this challenge, we first outline three types of object hallucinations frequently observed in the detailed captions: (1) Object Existence Hallucination - The detailed image description references objects that are not present; (2) Object Attribute Hallucination - The detailed image description inaccurately characterizes objects, misrepresenting attributes such as color, shape, and size; (3) Object Relationship Hallucination - The detailed image description inaccurately depicts the relationships or interactions among objects, including erroneous relative positions, interaction states, and actions involving two or more objects. In this work, we mainly focus on defining the metric, analyzing the cause, and addressing the the problem of object existence hallucination.
@@
Evaluating detailed captions is inherently complex. Some of the efforts, including benchmarks like POPE (Li et al., 2023e), evaluate object hallucination using VQA. Such a bias towards VQA-based evaluations might result in an incomplete assessment of detailed captions, which requires obtaining a comprehensive view of visual details. To bridge this gap, we introduce CCEval, designed specifically for object existence hallucination in detailed captions. To avoid the model gaining an unfair advantage by favoring shorter descriptions, CCEval maintains consistency in metrics such as average sentence length and the number of objects (see Sec. 2). Notably, even models that well-performed on VQA-based object hallucination benchmarks showed substantial hallucinations when evaluated with CCEval.
@@
In our exploration to uncover the underlying cause of object existence hallucination, we look into various factors including the size of the language decoder, the quantity, quality, and granularity of instruction data, and the input resolution to the vision encoder. We conclude the most crucial factor to be the alignment between objects mentioned in training caption and those vision encoder can perceive. During training of LMMs, the goal is to establish a one-to-one correspondence between objects mentioned in the caption and those present in the image. Objects successfully grounded by the vision encoder form accurate associations, internalizing them as contextual knowledge. Conversely, objects in the language that the vision encoder fails to ground create word-word semantic associations, which can be attributed to the generalization from the parametric knowledge within the model’s parameters. During inference, when the model draws from such parametric knowledge, any misalignment can manifest as hallucination, as the model attempts to “guess” details not grounded by the vision module.
@@
To address such hallucination, we are motivated by that not all hallucination is bad, and it is more desirable to control the generalization rather than outright removal of all imagined objects. Recognizing the significance of both contextual and parametric knowledge in ensuring generation reliability, we present HallE-Control, a novel approach to control the extent of expressed hallucination or parametric knowledge. We curate a 33k dataset similar to LLaVA (Liu et al., 2023d), incorporating both pure contextual knowledge and a blend of contextual knowledge with marked parametric knowledge. Leveraging this dataset, we train a lightweighted single linear layer to control over the frozen LMM. As demonstrated in Figure 1, a singular continuous parameter adjustment (e.g. ) during inference enables the model to produce detailed captions with only contextual knowledge (e.g., ) or blend with parametric knowledge (e.g., ). Furthermore, the inferred objects from parametric knowledge are automatically highlighted with distinct tokens (e.g., [object]) for human reference. This method offers the advantage of preserving object count and coverage as well as sentence length, while effectively control object existence hallucination.
@@
A comprehensive analysis on LMM components that influence hallucination, with a specific focus on alignment issues in the vision encoder and instruction data;
The first approach to control object existence hallucination within detailed captions;
A novel evaluation method for detailed caption object existence hallucination, with metrics such as object count, coverage, and average sentence length, alongside hallucination assessment.
@@ 2 Hallucination Analysis
Object existence hallucination can be influenced by several factors, including the language decoder, instruction data, and vision encoder. In our analysis, we address each factor individually. For a diverse methodological exploration, we select LLaVA, InstructBLIP (Dai et al., 2023), and Shikra (Chen et al., 2023): LLaVA and Shikra share the same model structure; Shikra and InstructBLIP use mixed-dataset and multi-task instruction data; InstructBLIP finetunes only Q-former, while the other finetune projector and LLM. More details about models can be found in Appendix.
@@ 2.1 Benchmarks
There are two primary methods, VQA-based and caption-based benchmarks, for evaluating object existence hallucination in LMMs.
@@
VQA-based benchmarks pose questions about objects within images. For a model to be considered hallucination-free, it should address these visual questions accurately. Notably, a large proportion of questions are simply binary, typically asking about the presence or attributes of objects.
The POPE benchmark evaluates object existence hallucination by a polling-based query method, consisting of a series of yes/no questions on sampled objects from visual instructions. POPE contains three sets: random, popular, and adversarial. These subsets respectively focus on randomly selected objects, frequently occurring objects, and those objects that co-occur in training sets. We choose POPE evaluation on MSCOCO (Lin et al., 2014) dataset. MME (Fu et al., 2023) coarse-grained recognition construct yes/no questions similarly but selects objects at random. This benchmark has 30 images, with each image paired with two questions: one positive and one negative.
In Table 1, LLaVA7B exhibits the greatest degree of hallucination, whereas Shikra outperforms other models in both POPE and MME. Specifically, Shikra shows a significantly higher F1 score in both POPE-popular and POPE-adversarial categories, while LLaVA7B displays the lowest. Additionally, Shikra’s ”Yes” ratio is closer to a balanced 50% compared to other models. However, in subsequent sections, we demonstrate that these observations from VQA-based benchmarks are not consistent with those from caption-based benchmarks.
@@
| CHAIRs | CHAIRi | Avg. Length | Avg. Object | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object | |
| LLaVA7B | 24.1 | 9.1 | 42.5 | 3.7 | 72.00 | 19.7 | 32.74 | 92.27 | 9.19 |
| LLaVA13B | 60.6 | 18.4 | 90.2 | 7.6 | 79.00 | 23.80 | 33.56 | 108.02 | 9.28 |
| Shikra7B | 59.1 | 16.6 | 91.2 | 7.5 | 83.00 | 24.40 | 33.29 | 109.37 | 9.10 |
| InstructBLIP7B | 1.4 | 1.7 | 2.3 | 0.8 | 72.00 | 22.30 | 29.76 | 108.42 | 8.04 |
@@
Table 2(left) reveals that while InstructBLIP exhibits minimal object existence hallucination, it averages a mere 0.8 objects per sentence. In contrast, LLaVA13B and Shikra manifest more hallucination, but they also generate more detailed captions, with as many as 7.6 and 7.5 objects per sentence, respectively. We find comparing object hallucinations is impractical when there is a significant disparity in average sentence length and number of objects.
Apart from these disparities, the use of a hard-coded ground truth set is another challenge. To counter these challenges, we introduce CCEval, a GPT-4 assisted evaluation for detailed captions. We first prompt LMMs to generate detailed captions on 100 randomly sampled images from Visual Genome (Krishna et al., 2017). Subsequently, utilizing GPT-4’s in-context learning capabilities, we extract individual objects from these captions and identify hallucinated ones by referencing the provided ground truth objects. On top of CHAIR Rohrbach et al. (2018), we introduce ”coverage” metric to ensure that the captions are detailed enough. This metric computes the ratio of objects in the caption that match the ground truth to the total number of ground truth objects. We additionally record and balance the average number of objects as well as the average length of captions across all cases. More details of CCEval can be found in Appendix.
As reflected in Table 2(right), when subjected to consistent constraints—average sentence length approximately 100 words and around 9 objects per sentence—all models exhibit comparably sub-optimal results. Interestingly, while Shikra surpass other models in VQA-based benchmarks, especially in the POPE , it under-performs in CCEval. This suggests that hallucination in detailed captions is not consistently captured by VQA-based evaluations.
@@
| POPE - Popular | LLaVA7B | 65.07 | 59.60 | 93.53 | 72.81 | 78.47 |
@@
| POPE - Adversarial | LLaVA7B | 57.07 | 54.07 | 93.93 | 68.63 | 86.87 |
@@
| Benchmark | Model | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
| CCEval (Ours) | LLaVA7B | 82.00 | 25.30 | 33.58 | 109.89 | 9.31 |
@@ 2.2 Language Decoder
We investigate if expanding the size of the language backbone can mitigate object existence hallucination. As detailed in Table 3, the language decoder of LLaVA is increased from 7B to 33B, and for InstructBLIP, it is increased from 7B to 13B. The result shows that hallucination for LLaVA reduced for POPE but not for CCEval. For InstructBLIP, and is reduced by 8 and 5.6 on CCEval, respectively. However, although there is a gain for scaling up language backbone, it is not consistent or salient from the observation, suggesting language decoder is not a primary factor in reducing hallucination.
@@ 2.3 Data
Similar to our approach with the language decoder, we begin by scaling up the volume of instruction finetuning data, ranging from 80K to 2.4M. As illustrated in Table 4, the LLaVA7B model, finetuned on 80K instruction data, exhibits fewer object existence hallucinations compared to the models finetuned on 150K and SVIT (Zhao et al., 2023a). The result suggests extra data without quality guarantee may increase hallucination for both VQA-based and caption-based evaluations. Given that all three datasets are generated by GPT-4, we question the quality of the data. LRV also raises this concern, suggesting that the training data itself might contain hallucinations. Some examples of training data are presented in the Appendix. Interestingly, we find that GPT-4 does not introduce additional object existence hallucination at all – the issue lies in MSCOCO ground truth objects that are given to GPT-4 and to be asked to strictly include in generated captions. These GT objects are challenging even for human observers to ground, due to factors like size, resolution, and occlusion. This led us to hypothesize that the vision encoder might also struggle to ground these objects effectively.
@@
| Benchmark | Finetune Data | CHAIRi | CHAIRs | Coverage | Avg. Length | Avg. Object |
@@
| Benchmark | Vision Encoder | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
@@
| Benchmark | Vision Encoder | CHAIRs | CHAIRi | Coverage | Avg. Length | Avg. Object |
@@ 2.4 Vision Encoder
Intuitively, increasing image resolution enhances model’s perception of finer details, thus making the grounding of objects mentioned in the caption easier. To verify our hypothesis from the previous section, we increase the input image resolution for the vision encoder. Specifically, for our evaluation, the resolution for LLaVA7B was incremented from 224x to full resolution using a sliding window approach for efficiency, as detailed in the Appendix. Table 6 shows a constant decrease in hallucination and increase in object coverage. Additionally, we assess LLaVA with Llama 213B, varying the resolution from 112x to 336x. For the 112x112 resolution, the original image was downscaled to 112x and subsequently upscaled to 224x before being input to CLIP-Large-224x. Table 5 gives a consistent observation that larger input resolution can reduce hallucination in image captions.
Conclusion. Through our systematic analysis of object existence hallucination, we summarize several insights: (1) Enlarging the language decoder does mitigate hallucination, but improvements are not huge. (2) Expanding the volume of instruction data actually increases hallucination. Upon inspection of training data, we find certain objects described in the captions might not be grounded by the vision encoder. (3) To validate our hypothesis in (2), we show improving input image resolution significantly reduces hallucination by enhancing model grounding ability.
Reflecting on these findings, we attempt to provide an explanation for the reason of object existence hallucination in detailed captions. The process of image captioning in LMMs can be perceived as a form of information mapping or translation. Ideally, the goal is to have a direct one-to-one correspondence between objects identified in the image and those mentioned in the captions. Objects successfully grounded by the vision encoder form accurate correspondence, making this as contextual knowledge in the model, following (Neeman et al., 2023). When objects in training caption fail to ground by the vision encoder, the model learns parametric knowledge, the knowledge encoded in the model’s parameter. This kind of knowledge is the association of objects in the language with other words instead of with corresponding image object feature. During inference, when the model draws from parametric knowledge, it attempts to “guess” details not grounded by the vision module and is perceived as object existence hallucination.